# Impact of flavour coupling on $SO(10)$-inspired leptogenesis

### v1

@@ 5 Final remarks
We could benefit of the analytical description of SO10INLEP given in [33] to re-derive the allowed region in the space of low energy neutrino parameters including flavour coupling. The results confirm the broad picture already discussed in previous papers but we noticed how we are currently entering a new stage for the prospect of testing SO10INLEP. There are also interesting new aspects introduced by flavour coupling. Let us briefly summarise our results.
@@
  • The cosmological upper bound Eq. (2) definitively rules out quasi-degenerate neutrinos and the latest upper bounds from $0\nu\beta\beta$ and tritium decay experiments confirm this so far. Absolute neutrino mass scale experiments are then starting to test the most interesting absolute neutrino mass range for the opportunities to test SO10INLEP, since for $10\,{\rm meV}\lesssim m_{1}\lesssim 30\,{\rm meV}$ there is a stringent upper bound on the atmospheric mixing angle that excludes second octant.
@@
  • An interesting new feature introduced by flavour coupling is that new muon-dominated solutions appear in a portion of the space of parameters that would be otherwise excluded if flavour coupling is not taken into account. There are also some completely new electron-dominated solutions appearing (in already allowed region of parameter space) but these are quite marginal and do not seem to have a particular interest.
@@
  • We have confirmed the existence of ST-SO10INLEP solutions even when flavour coupling is taken into account. The allowed region gets even larger and, importantly, the lower bound on the absolute neutrino mass scale gets more relaxed. The washout of a large initial pre-existing asymmetry $N^{\rm p,i}_{B-L}\sim 10^{-3}$ requires $m_{ee}(m_{1})\gtrsim 4(3)\,{\rm meV}$. This is still large enough that future generation $0\nu\beta\beta$ experiments will be able to test it. The fact that ST-SO10INLEP is protected by flavour coupling is somehow a built-in feature: since $K_{1\tau}\lesssim 1$, this necessarily limits the possibility to have a too large leak of the pre-existing asymmetry from the $e$ and $\mu$ flavours into the $\tau$ flavour so that the pre-existing asymmetry can be still efficiently washed-out for a subset of the SO10INLEP solutions. This is an important result of our paper since it shows that the ST-SO10INLEP scenario is stable under perturbations introduced by a more sofisticate calculation of the asymmetry.
@@
  • Our procedure, providing an approach toward a map of $SO(10)$-inspired models, can also be useful within the quest for a realistic grandunified model since it gives a complementary analytical insight and a global picture of the solutions. Indeed it has provided quite a useful insight in connection with a recent identification of a realistic fit of fermion parameters within a minimal $SO(10)$ scenario that is also able to reproduce the observed baryon asymmetry with $N_{2}$-leptogenesis [56]. This falls within the category of SO1OINLEP scenarios but extended to allow for a large, in fact maximal, value of $\theta_{23}^{L}$. [^13] Recently, it was also emphasized the role that a triplet Higgs might have in $SO(10)$-leptogenesis and in the search of realistic minimal $SO(10)$-models [57]. For a recent mini-review on studies embedding high scale thermal leptogenesis within grandunified models see [58].
@@
In conclusion, our results represent the first step toward a systematic understanding of theoretical uncertainties within SO10INLEP that is particularly important in view of future experimental expected progress both in long baseline neutrino oscillation experiments and absolute neutrino mass scale experiments. It is a very exciting time for the possibility of testing SO10INLEP. Both long baseline and absolute neutrino mass scale experiments will gradually test its predictions and constraints. In particular, $0\nu\beta\beta$ experiments are finally entering the mass range that is predicted by the strong thermal version of SO10INLEP. An account of flavour coupling has confirmed and even strengthened the picture, making the predictions more robust. At the same time, it also provides useful information for the optimization of successful strategy toward the identification of a grandunified realistic model.

### v2

@@ 5 Final remarks
We could benefit of the analytical description of SO10INLEP given in full to re-derive the allowed region in the space of low energy neutrino parameters including flavour coupling. The results confirm the broad picture already discussed in previous papers but there are also interesting new aspects introduced by flavour coupling. Moreover, in general, we also noticed how we are currently entering a new stage for the prospect of testing SO10INLEP. Let us briefly summarise our main results.
@@
  • The cosmological upper bound Eq. (2) definitively rules out quasi-degenerate neutrinos and the latest upper bounds from $0\nu\beta\beta$ and tritium decay experiments confirm this so far. Absolute neutrino mass scale experiments are then starting to enter the most interesting absolute neutrino mass range for the possibility of testing SO10INLEP, since for $10\,{\rm meV}\lesssim m_{1}\lesssim 30\,{\rm meV}$ there is a stringent upper bound on the atmospheric mixing angle that excludes second octant.
@@
  • An interesting new feature introduced by flavour coupling is that new muon-dominated solutions appear in a portion of the space of parameters that would be otherwise excluded if flavour coupling is not taken into account. There are also some completely new electron-dominated solutions appearing (inside an already allowed region of parameter space) but these are quite marginal and do not seem to have a particular interest.
@@
  • We have confirmed the existence of ST-SO10INLEP solutions even when flavour coupling is taken into account. The allowed region gets even larger and, importantly, the lower bound on the absolute neutrino mass scale gets more relaxed. The washout of a large initial pre-existing asymmetry $N^{\rm p,i}_{B-L}\sim 10^{-3}$ requires $m_{ee}(m_{1})\gtrsim 4(3)\,{\rm meV}$. This is still large enough that future generation $0\nu\beta\beta$ experiments will be able to test it. The fact that ST-SO10INLEP is protected by flavour coupling is somehow a built-in feature: since $K_{1\tau}\lesssim 1$, this necessarily limits the possibility to have a too large leak of the pre-existing asymmetry from the $e$ and $\mu$ flavours into the $\tau$ flavour so that the pre-existing asymmetry can be still efficiently washed-out for a subset of the SO10INLEP solutions. This is an important result of our paper since it shows that the ST-SO10INLEP scenario is stable under perturbations introduced by a more sophisticate calculation of the asymmetry.
@@
  • Our procedure, providing an approach toward a map of $SO(10)$-inspired models, can also be useful within the quest for a realistic grandunified model since it gives a complementary analytical insight and a global picture of the solutions. Indeed it has provided quite a useful insight in connection with a recent identification of a realistic fit of fermion parameters within a minimal $SO(10)$ scenario that is also able to reproduce the observed baryon asymmetry with $N_{2}$-leptogenesis Babu:2024ahk . This falls within the category of SO1OINLEP scenarios but extended to allow for a large, in fact maximal, value of $\theta_{23}^{L}$. [^18] Recently, it was also emphasized the role that a triplet Higgs might have in $SO(10)$-leptogenesis and in the search of realistic minimal $SO(10)$-models Fong:2025aya . For a recent mini-review on studies embedding high scale thermal leptogenesis within grandunified models see Malinsky:2025jwg .
@@
In conclusion, our results represent the first step toward a systematic understanding of theoretical uncertainties within SO10INLEP that is particularly important in view of future experimental expected progress both in long baseline neutrino oscillation experiments and absolute neutrino mass scale experiments. It is a very exciting time for the possibility of testing SO10INLEP. Both long baseline and absolute neutrino mass scale experiments will gradually test its predictions and constraints. In particular, $0\nu\beta\beta$ experiments are finally entering the mass range that is predicted by the strong thermal version of SO10INLEP. An account of flavour coupling has confirmed and even strengthened the picture, making the predictions more robust. At the same time, it also provides useful information for the optimization of a successful strategy toward the identification of a grandunified realistic model.